OpenStudy (anonymous):

Convert 510° from degrees to radians. Use the value of π found on a calculator and round answers to four decimal places, as needed.

OpenStudy (anonymous):

About 8.9012.

OpenStudy (anonymous):

@saze what about in pi form??

OpenStudy (anonymous):

I'm not certain what you mean by pi form, but the equation I used to get this was \[r = d \times \frac{ \pi }{ 180 }\] where r is number in radians, and d is number in degrees.

OpenStudy (anonymous):

@saze so is it 17pi/3?

OpenStudy (anonymous):

That gives us 17.8; I'd just leave it in \[\frac{ 510\pi }{ 180 }\]

OpenStudy (anonymous):

thankyou!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

If you simplify it, it's 17pi over six.